OpenStudy (fantageplayer):

Consider the functions f(x) = 2x + 1 and g(x) = x2 − 10. What is the value of f[g(3)]? 39 −3 −7 −1

OpenStudy (faiqraees):

First find the value of the function g(3)

OpenStudy (fantageplayer):

3x2−30?

OpenStudy (faiqraees):

No input 3 instead of x

OpenStudy (fantageplayer):

oh so divide then?

OpenStudy (faiqraees):

no g(3)=(3)²-10

OpenStudy (fantageplayer):

but dont you divide (3)^2-10/3 if you want g?

OpenStudy (fantageplayer):

sorry im bad at this

OpenStudy (faiqraees):

No no it is not 3g it is g(3) which is a different thing. g(3) just means the value when x is 3

OpenStudy (fantageplayer):

ohhhhhhhhhhhhhhhhh

OpenStudy (faiqraees):

for example if x = 1 then x²-10 = -9 so g(1) = -9

OpenStudy (fantageplayer):

i see

OpenStudy (fantageplayer):

does x = 3 for x too?

OpenStudy (fantageplayer):

@radar @FaiqRaees

OpenStudy (fantageplayer):

does x = 3 for f too?*

OpenStudy (faiqraees):

no did you find the answer for g(3)

OpenStudy (fantageplayer):

how do i do that?

OpenStudy (faiqraees):

for g(3) replace 3 in place of x

OpenStudy (fantageplayer):

so thenit would be -3?

OpenStudy (faiqraees):

no g(x) = x²-10 g(3) = 3²-10 g(3)=9-10

OpenStudy (radar):

FaiqRaees is doing an excellent job, focus, study the posts.

OpenStudy (mrcoolguy):

sorry

OpenStudy (fantageplayer):

ohhhh i mean -1

OpenStudy (faiqraees):

yep since g(3) = -1 so you can replace that into f(g(3)) f(g(3)) = f(-1)

OpenStudy (faiqraees):

now find the value of f(-1)

OpenStudy (fantageplayer):

-1?